OpenStudy (anonymous):

compute the limit lim (8x-x^2+3)^3 as x goes to 2

OpenStudy (across):

\[\lim_{x\to2}(-x^2+8x+3)^3=(-(2)^2+8(2)+3)^3=?\]

OpenStudy (anonymous):

OK SO I JUST SOLVE RIGHT?

OpenStudy (across):

Yes

OpenStudy (across):

However, when you start seeing things like\[\lim_{x\to2}\frac{x^2+3x+2}{x-2},\]then you have to do some algebraic manipulation.

OpenStudy (anonymous):

OK. FOR THE EQUATION ABOVE WOULD IT (-(4)+16+9 ?

OpenStudy (across):

Where did you get that \(9\) from?

OpenStudy (anonymous):

I got it from the last part of the equation where it says 3)^3. Or did I do that wrong

OpenStudy (turingtest):

yes, that's wrong ^3 is the last step; do what's in the parentheses first

OpenStudy (turingtest):

and ^3 does not mean times 3, it means to the third power

OpenStudy (anonymous):

so would it be (-4+16+3)^3

OpenStudy (across):

Yes. What does that evaluate to?

OpenStudy (anonymous):

-4+16+3= 15 so then I do 15^3 Right? so that would equal 3375

OpenStudy (across):

That's right.

OpenStudy (anonymous):

Thank you for your help